OpenStudy (anonymous):

need help solving for the variable in the equations. will post the equation below. and will give metal to best answer.

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 } d = \frac{ 5 }{ }\]

OpenStudy (anonymous):

opps, sorry. it is \[\frac{ 5 }{ 8 }\]